OpenStudy (anonymous):

What are the x-intercepts of the parabola x2 − 3x − 28?

OpenStudy (anonymous):

Answers....

OpenStudy (anonymous):

x^2 - 3x - 28 = 0 (x + 3)(x - 7) = 0 x = -3, x = 7

OpenStudy (anonymous):

wait oops, typo messed me up.

OpenStudy (anonymous):

x^2 - 3x - 28 = 0 (x + 4)(x - 7) = 0 x = -4, x = 7 So the answer is the last.

OpenStudy (anonymous):

Are you sure about that ? Because I don't have any of those options.. @melissa780

OpenStudy (anonymous):

I corrected myself, do you see it now? Sorry about that. :x

OpenStudy (anonymous):

yeah I saw that? But I don't have any of that for an option to choose....

OpenStudy (anonymous):

Yes....... you do. Look at your last option. I literally wrote it out for you, love. x = -4, x = 7 Which translates to (-4, 0) and (7, 0) which as I already said, it's your last option in the exam.

OpenStudy (anonymous):

OH MY GOSH. Whoops ! Sorry. Psh I knew that O_O

OpenStudy (anonymous):

Don't worry about it, lovely!